# Time-Series Forecasting and Sequence Learning Using Memristor-based Reservoir System

### v1

@@ 3. System Design and Implementation
It is important to highlight that each crossbar in the reservoir layer uses a 1M configuration. Although this configuration is well-known to possess 10$\times$ smaller footprint as compared to 1T1M (Yakopcic, 2014), it is harder to configure. The reconfigurability is an essential feature in ESN hardware implementation as it endows the network with an additional degree of freedom to continuously change its structure and sparsity level, ensuring network stability. In this work, the reconfigurability (synaptogenesis) is enabled through the use of the Ziksa writing scheme, which we previously proposed in (Zyarah et al., 2017). Ziksa is used to suppress the effect of the synaptic connections that are not involved in computation, promoting network sparsity. The suppression occurs by setting the conductance of the memristor devices emulating the positive portion of the weights ($M^{+}$) to be equal to that representing the negative portion of the weights ($M^{-}$), such that the net current flowing in these synapses for a given input is zero. Our approach significantly reduces the resources required to enable sparsity unlike the work in (Kume et al., 2020), which suggests using switching transistors and 1-bit memory cells or flip-flops to control each MOSFET device in the crossbar.
@@ 3.1. ESN Accelerator Training
As discussed in the overview section, the training in ESNs is confined solely to the readout layer, and it can either be performed analytically via ridge regression or iteratively by using LMS. In this work, we choose LMS to ensure fast learning and to minimize resource usage. However, the LMS training equation consists of two terms: gradient and regularization. The gradient is first estimated via computing network output error, the difference between the predicted value and the ground truth which presented by the training circuitry block. Typically, a subtractor, a non-inverting operational amplifier, is used for this purpose. The output of the subtractor is multiplied by the activation of the reservoir neuron to estimate the gradient required to modulate the corresponding memristor conductance. The gradient is digitized by using a 6-bit Flash-ADC (optimized for power constrained devices) and the output is sent to a global controller where it gets temporally stored. Then, the regularization term is estimated as follows. First, a fixed test voltage [^2] Due to technology limitations, the test signal ($V_{test}$) is set to be 50mV to avoid any undesired clipping or distortion in the generated output. is applied to the memristor required to be updated ($M_{u}$). Then, the output voltage will be captured again by the ADC and the result will be sent to the global controller to estimate the value of the memristor conductance according to (4).
@@
Once the memristor conductance is estimated, is it multiplied by the regularization factor and the result will be added up to the gradient in order to estimate the amount of change that needs to be made to a memristor. This amount ($\Phi$) will be translated into a pulse signal of fixed amplitude and variable duration (duration reflects the amount of change to be made to a memristor) using the conversion circuit shown in Fig. 5, in which the output ($V_{m}$) is given by (5) and (6). Here, $V_{int}$ denotes the output of the integrator, and $V_{r}$ is a reference input voltage, which along with R and C controls the speed of integration.
@@
(6) $V_{m}=\begin{cases}T_{h},&\Phi>V_{int}\\ 0,&otherwise\end{cases}$
It is imperative to mention that the change of memristors occurs in a sequential fashion, one memristor at a time. This is to make more precise tuning and to reduce the power consumption during the training process. While such a training approach is known to be slow, we speed it up by means of gradient sparsification, which limits the changes to the devices with gradient values above a predefined threshold ($\Theta$). It is important to highlight that the training operation is verified in Cadence Virtuoso for a small-scale network (1x20x1), whereas the large-scale ESN verification is done using a python model. The python model takes into consideration all the hardware constraints, such as limited range and precision of the weight representation, non-linearity of the memristor devices, leakage of the sample and hold circuit, etc.
@@ 4.1. Design Space Exploration
In order to select the best hyperparameters that result in a compact design and best performance, the particle swarm optimization algorithm (PSO) is used. The algorithm is integrated with the ESN software model and set into constrained mode, in which the feasible points for hyperparameters are defined based on the hardware constraints. This includes the limited range of weight representation, crossbar size, precision, etc. In this work, the PSO employs a swarm of 50 particles, which are chosen to optimize for the following: the number of reservoir neurons that controls the high-dimensional space of feature extraction, frequency of synaptic updates ($n_{up}$), sparsity level within the reservoir layer which influences the computational cost, leakage rate to moderate the speed of the reservoir dynamic updates, regularization term to mitigate the overfitting, and learning rate to regulate the convergence speed. The algorithms run for 50 epochs with the target of minimizing the wMAPE of real-world benchmarks when performing forecasting tasks.
@@ 4.2. Device Non-idealities and Process Limitations
There are several operation and reliability concerns when using memristor devices to model the synaptic connections in neural network accelerators, such as endurance limitation and device non-idealities. The endurance limitation is highly impacted by several factors, including the programming voltage, device structure, and device material. For instance, using excessive writing voltage can lead to a major deterioration in memristor endurance, whereas the proper use of programming voltage leads to a significant improvement (Wang et al., 2015). However, crossing the endurance limits may drive the memristor devices to exhibit various behaviors which may severely impact network performance. For example, the authors in (Yang et al., 2010) observe a reduction in the metal-oxide memristor (Pt/TaOx/Ta) resistance ratio once the switching cycles surpass $6\times 10^{9}$. Once the device switching window is collapsed, the device will be stuck at the reset resistance rather than being shorted. In (Kim et al., 2016), the authors observed different behavior. After surpassing the device switching cycles limit, the memristor experienced a stuck-at failure. In this work, since we are modeling the physical characteristics of the device proposed by (Yang et al., 2010), we will consider that the memristors experience stuck-at failure when passing the endurance limit. The endurance limit considered is $1\times 10^{9}$ cycles when tuned with a fixed voltage pulse of $\pm$1.2v for set and reset.
When it comes to the device non-idealities, we have cycle-to-cycle and device-to-device variabilities, which characterize the time-varying stability of memristors and their uniformity when integrated into crossbar structure (Vourkas and Sirakoulis, 2016). In our work, variabilities due to device material and imperfect manufacturing process (Amat et al., 2016), have been considered. The cycle-to-cycle, which is the write variation to memristor devices, is modeled by adding noise to the learning rule. The device-to-device variability is confined to the CMOS devices and memristor resistance range and threshold. Regarding the process variation and limitations, it is considered for the developed 6-bit flash ADC, comparators, and operations amplifiers [^3] All analog units are tested for fabrication process, voltage supply, and ambient temperature variations using corner analysis..
Figure 6. Fitting of the used memristor device model to the physical device characteristics.
@@ 4.3. Device Model
Table 1. The memristor device parameters used in the reservoir and readout layers.
Parameter Value [Reservoir] Value [Readout]
@@
Full switching pulses 41 41
@@ 4.4. Verification Benchmark
In order to verify the operation of the proposed memristor-based ESN accelerator several univariate benchmarks have been used in the time-series forecasting task:
  • PJM Energy: this dataset holds 145366 samples representing regional energy consumption in the USA. The energy consumption is recorded every hour for the period 01/01/2010 to 07/01/2018. ESN is employed for short-term energy consumption forecasting, i.e. the energy consumption of the next 50-100 hours.
  • Daily temperature: daily minimum temperature in Melbourne, Australia recorded between the period (1981-1990). The dataset contains 3,605 noisy samples smoothed with moving window average (window size=5) to reduce the noise effect. The ESN is used to predict the temperature for the next day.
  • Mackey-Glass: dataset generated by non-linear time-delayed differential equation modeling a chaotic system (Mackey and Glass, 1977), see (10). This dataset is widely used as a benchmark for time-series forecasting tasks. In this work, the parameters of the Mackey-glass equation are set as follows: $\beta=0.25,\gamma=0.1,\tau=18,\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ n=10$, whereas the number of generated examples used during the forecasting task is 4000.
@@
  • NARMA10: The classical non-linear autoregressive moving average (NARMA) introduced by (Atiya and Parlos, 2000). It represents a dynamical system that is difficult to model due to the non-linearity and long-term dependencies. In this work, 4000 samples from the $10^{th}$ order NARMA system are employed for the forecasting task. The system is defined by (11), where $y(t)$ and $s(t)$, respectively, are the output and the input of the system at time $t$.
@@
All the benchmark features are scaled to be ranged between 0 and 1 so that they have the same interval as the reservoir and output neurons. Furthermore, the first 100 samples are dedicated to the initial washout period.
@@ 5.1. Time-series Forecasting
In order to quantify the performance of the proposed ESN memristor-based accelerator, the time-series forecasting is evaluated using the weighted mean absolute percentage error (wMAPE) metric, which is given by (12). With weights ($\varpi_{i}$) set to unity, the wMAPE can determine the average difference between the actual and predicted values while capturing the substantial fluctuations in magnitude (Miraftabzadeh et al., 2023).
@@
Fig. 7-(a) depicts the wMAPE of the software and the developed hardware model of the ESN network when predicting the energy consumption of the load for the next 50 hours. It can be noticed that the wMAPE, which is recorded every time 250 samples are introduced to the network, starts with a high value and then gradually degrades over time as the network learns and captures short and long-term dependencies. With the absence of the regularization term in the learning equation seems that the network takes a longer time to adapt. Furthermore, the network struggles to re-adapt when there is a major change in input patterns. When it comes to the developed hardware model, which uses regularization, a similar trend as the software model has been observed. However, one may see there is a gap in performance ($\sim$4.8%) between the software and hardware model, which is attributed to several reasons. Among these, are the non-idealities of memristor devices, limited precision of the ADC used when modulating the conductance of the memristors, undesired leakage in the sample and hold circuit leading to a drop in the stored charges.
Table 2. Forecasting wMAPE of uni-variate stationary and non-stationary benchmarks using the proposed memristor-based ESN when trained with LMS, LMS with L2 regularization, and using point and LI neurons in the reservoir layer.
Benchmark Forecast LMS (LIF) LMS+L2 LMS+L2 (LIF)
PJM-Energy 50-Step 0.093$\pm$0.0073 0.107$\pm$ 0.012 0.08$\pm$0.001
100-Step 0.097$\pm$0.0093 0.120$\pm$0.019 0.085 $\pm$0.0073
Macky-Glass 50-Step 0.082$\pm$0.0003 0.237$\pm$0.0091 0.079$\pm$0.0001
100-step 0.087$\pm$0.0005 0.236$\pm$0.0093 0.081$\pm$0.00021
Daily-Temp 50-Step 0.104$\pm$0.0013 0.259$\pm$0.01 0.108$\pm$0.0011
100-step 0.103$\pm$0.005 0.261$\pm$0.011 0.109$\pm$0.0027
NARMA10 50-Step 0.197$\pm$0.039 0.198$\pm$0.058 0.191$\pm$0.077
100-Step 0.212$\pm$0.087 0.231$\pm$0.0027 0.184$\pm$0.039
Table 2 shows the wMAPE when forecasting data from stationary and non-stationary benchmarks for various time steps. The wMAPE is calculated when training the proposed memristor-based ESN accelerator with LMS, and LMS with L2 regularization using point and leaky-integrated neurons in the reservoir layer. It can be seen that having leaky-integrated neurons significantly improves network performance as it enables ESN to form temporal dependencies with controllable dynamic updates.
@@ 5.2. Device Failure Effect
There are several types of device failure a network has to deal with especially when using memristor devices. The memristor devices may experience a deviation in device characteristics such as a change in the device conductance range due to excessive device switching, known as an aging fault, or it may have a fault in the device right after the fabrication process such as stuck-at-fault. In this work, we will mainly focus on the stuck-at-fault because it is most common and may severely impact network performance (Xu et al., 2018). There are three types of stuck-at fault: stuck-at, stuck-on, and stuck-off. Previous work provided that stuck-at replicates the memristor device failure after the forming process and such fault has marginal effect on network performance (Zyarah and Kudithipudi, 2019b). Thus, this work solely investigates the impact of stuck-on (high-conductance state) and stuck-off (low-conductance state) faults on predicting future events in time-series data. In Fig. 7, we show the change in wMAPE when predicting the energy consumption of the load for the next 50 hours in the presence of various levels of stuck-on and stuck-off faults occurring in the readout and reservoir layers. The fault impact is investigated in both 1M1R and 2M crossbar structure. In 1M1R structure, it is found that regardless of the fault type, below 10% its impact can be deemed marginal causing $\pm$1.3change in the wMAPE value. However, this impact exacerbates beyond 10% and seems to have an exponential negative impact on wMAPE, see Fig. 7-(b). In 2M structure, we may observe similar impact, but it can be suppressed by leveraging the programmable nature of memristor devices. When using two memristors to represent each synaptic weight, one can: i) leverage the intact memristors to tune the weights modeled by faulty devices and here the weights will not be frozen but rather have limited dynamic range, ii) enforce sparsity, i.e. zero weight values, via equating the conductance of the intact and faulty devices. Fig. 7-(c) illustrates the wMAPE in the presence of stuck-on and stuck-at faults when we enforce sparsity.
Figure 7. (a) The wMAPE of the developed ESN models (SW and HW) while predicting the load energy consumption for 50 hours computed every 250 samples. (b) and (c) The impact of various levels of the stuck-on and stuck-off fault in memristor devices on the network performance while performing time-series forecasting task when using 1M1R and 2M crossbar structures, respectively.

### v2

@@ 3. System Design and Implementation
It is important to highlight that each crossbar in the reservoir layer uses a 1M configuration. Although this configuration is well-known to possess a 10$\times$ smaller footprint as compared to 1T1M (Yakopcic, 2014), it is harder to configure. Reconfigurability is an essential feature in ESN hardware implementation as it endows the network with an additional degree of freedom to continuously change its structure and sparsity level, ensuring network stability. In this work, reconfigurability (synaptogenesis) is enabled through the use of the Ziksa writing scheme, which we previously proposed in (Zyarah et al., 2017). Ziksa is used to suppress the effect of the synaptic connections that are not involved in computation, promoting network sparsity. This suppression occurs by setting the conductance of the memristor devices emulating the positive portion of the weights ($M^{+}$) to be equal to that representing the negative portion of the weights ($M^{-}$), such that the net current flowing through these synapses for a given input is zero. Our approach significantly reduces the resources required to enable sparsity, unlike the work in (Kume et al., 2020), which suggests using switching transistors and 1-bit memory cells or flip-flops to control each MOSFET device in the crossbar.
@@ 3.1. ESN Accelerator Training
As discussed in the overview section, training in ESNs is confined solely to the readout layer, and it can be performed either analytically via ridge regression or iteratively by using LMS. In this work, we choose LMS to ensure fast learning and minimize resource usage. However, the LMS training equation consists of two terms: gradient and regularization. The gradient is first estimated by computing network output error, the difference between the predicted value and the ground truth as presented by the training circuitry block. Typically, a subtractor, a non-inverting operational amplifier, is used for this purpose. The output of the subtractor is multiplied by the activations of the reservoir neurons to estimate the gradient required to modulate the corresponding memristor conductance. The gradient is digitized using a 6-bit Flash-ADC (optimized for power constrained devices) and the output is sent to a global controller where it gets temporarily stored. Then, the regularization term is estimated as follows: First, a fixed test voltage [^3] Due to technology limitations, the test signal ($V_{test}$) is set to be 50mV to avoid any undesired clipping or distortion in the generated output. is applied to the memristor that needs to be updated ($M_{u}$). The output voltage is then captured again by the ADC, and the result will be sent to the global controller to estimate the value of the memristor conductance according to (4).
@@
Once the memristor conductance is estimated, it is multiplied by the regularization factor, and the result is added to the gradient in order to estimate the amount of change that needs to be made to a memristor. This amount ($\Phi$) is then translated into a pulse signal with a fixed amplitude and variable duration (duration reflects the amount of change to be made to a memristor) using the conversion circuit shown in Fig. 3, where the output ($V_{m}$) is given by (5) and (LABEL:integ1). Here, $V_{int}$ denotes the output of the integrator, and $V_{r}$ is a reference input voltage, which along with R and C, controls the speed of integration.
(5) $V_{m}=\begin{cases}T_{h},&\Phi>V_{int}\\ 0,&otherwise\end{cases}$
@@
It is imperative to mention that the changes in memristors occur in a sequential fashion, one memristor at a time. This approach allows for more precise tuning and reduces power consumption during the training process. While such a training approach is known to be slow, we speed it up by means of gradient sparsification, which limits the changes to the devices with gradient values above a predefined threshold ($\Theta$). The training operation is verified in Cadence Virtuoso for a small-scale network (1x20x1) [^4] The training is verified while considering memristor cycle-to-cycle and device-to-device variabilities. Variabilities with a normal distribution has been considered, with a mean defined by device parameters and a standard deviation equal to 10% of the mean., whereas the large-scale ESN verification is done using a Python model. The Python model takes into consideration all the hardware constraints, such as limited range and precision of the weight representation, non-linearity of the memristor devices, leakage in the sample and hold circuit, etc.
@@ 4.1. Design Space Exploration
In order to select the best hyperparameters that result in a compact design and optimal performance, the particle swarm optimization algorithm (PSO) is used. The algorithm is integrated with the ESN software model and set into constrained mode, in which the feasible points for hyperparameters are defined based on the hardware constraints. This includes the limited range of weight representation, crossbar size, precision, etc. In this work, the PSO employs a swarm of 50 particles, which are chosen to optimize for the following: the number of reservoir neurons that controls the high-dimensional space of feature extraction, frequency of synaptic updates ($n_{up}$), sparsity level within the reservoir layer which influences the computational cost, leakage rate to moderate the speed of the reservoir dynamic updates, regularization term to mitigate the overfitting, and learning rate to regulate convergence speed [^5] The leakage rate, regularization term, and learning rate are set to be global parameters.. The algorithm runs for 30 iterations with the target of minimizing the wMAPE of real-world benchmarks when performing forecasting tasks.
@@ 4.2. Device Non-idealities and Process Limitations
There are several operational and reliability concerns when using memristor devices to model the synaptic connections in neural network accelerators, such as endurance limitations and device non-idealities. The endurance limitation is highly impacted by several factors, including the programming voltage, device structure, and device material. For instance, using excessive writing voltage can lead to a major deterioration in memristor endurance, whereas the proper use of programming voltage leads to a significant improvement (Wang et al., 2015). However, crossing the endurance limits may drive the memristor devices to exhibit various behaviors which may severely impact network performance. For example, the authors in (Yang et al., 2010) observed a reduction in the metal-oxide memristor (Pt/TaOx/Ta) resistance ratio once the switching cycles surpass $6\times 10^{9}$. Once the device switching window is collapsed, the device will be stuck at the reset resistance rather than being shorted. In (Kim et al., 2016), the authors observed different behavior. After surpassing the device switching cycles limit, the memristor experienced a stuck-at failure. In this work, since we are modeling the physical characteristics of the device proposed by (Yang et al., 2010), we will consider that the memristors experience stuck-at failure when passing the endurance limit. The endurance limit considered is $1\times 10^{9}$ cycles when tuned with a fixed voltage pulse of $\pm$1.2v for set and reset.
When it comes to the device non-idealities, we have cycle-to-cycle and device-to-device variabilities, which characterize the time-varying stability of memristors and their uniformity when integrated into a crossbar structure (Vourkas and Sirakoulis, 2016). These non-idealities typically occur due to device material and imperfect manufacturing processes (Amat et al., 2016). However, in this work, the cycle-to-cycle variability of the device resistance range has been considered and is modeled as a variation in the weight range. For device-to-device variation, it is applied to the memristor threshold and resistance (write variation). Here, the write variation refers to the variability in the rate of change in device resistance during the learning process, which is modeled by adding noise to the learning rule. Regarding the process variation and limitations, it is considered for the developed 6-bit flash ADC, comparators, and operational amplifiers [^6] All analog units are tested for fabrication process, voltage supply, and ambient temperature variations using corner analysis..
@@ 4.3. Device Model
(7) $G_{mem}=\frac{w}{D}\times G_{on}+(1-\frac{w}{D})\times G_{off}$
Figure 4. Fitting of the used memristor device model to the physical device characteristics provided in (Yang et al., 2010)
Table 1. The memristor device parameters used in the developed ESN accelerator.
Parameter Value [Reservoir and Readout] Value [Leakage Cell]
@@
Full switching pulses 41 67
@@ 4.4. Verification Benchmark
In order to verify the operation of the proposed memristor-based ESN accelerator, several univariate benchmarks have been used in the time-series forecasting task:
  • PJM Energy: This dataset holds 145,366 samples representing regional energy consumption in the USA (PJM Interconnection, LLC, [n. d.]). The energy consumption is recorded every hour for the period 01/01/2010 to 07/01/2018. The ESN is employed for short-term energy consumption forecasting, i.e., the energy consumption for the next 50-100 hours.
  • Daily temperature: The daily minimum temperature in Melbourne, Australia recorded between the period (1981-1990) (Hyndman and Yang, 2018). The dataset contains 3,605 noisy samples. We smoothed the data with moving average (window size=5) to reduce the noise effect and used ESN to predict the temperature for the next day.
  • Mackey-Glass: This dataset is generated by non-linear time-delayed differential equation modeling a chaotic system (Mackey and Glass, 1977), see (10). It is widely used as a benchmark for time-series forecasting tasks. In this work, the parameters of the Mackey-glass equation are set as follows: $\beta=0.25,\gamma=0.1,\tau=18,\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ n=10$, whereas the number of the generated examples used in forecasting task with ESN is 4000.
@@
  • NARMA10: The classical non-linear autoregressive moving average (NARMA) was introduced by (Atiya and Parlos, 2000). It represents a dynamical system that is difficult to model due to its non-linearity and long-term dependencies. In this work, 4000 samples from the $10^{th}$ order NARMA system are employed for the forecasting task. The system is defined by (11), where $y(t)$ and $s(t)$, respectively, are the output and the input of the system at time $t$.
@@
All the benchmark features are scaled to range between 0 and 1 so that they have the same interval as the reservoir and output neurons. Furthermore, the first 100 samples are allocated to the initial washout period.
@@ 5.1. Time-series Forecasting
In order to quantify the performance of the proposed ESN memristor-based accelerator, time-series forecasting is evaluated using the weighted mean absolute percentage error (wMAPE) metric, which is given by (12). With weights ($\varpi_{i}$) set to unity, the wMAPE can determine the average difference between the actual and predicted values while capturing the substantial fluctuations in magnitude (Miraftabzadeh et al., 2023).
@@
Fig. 5-(a) depicts the wMAPE of the software and the developed hardware model of the ESN when using the PJM energy dataset to predict the energy consumption of the load for the next 50 hours. It is evident that the wMAPE, which is recorded every time 250 samples are introduced to the network, starts with a high value and then gradually degrades over time as the network learns and captures short and long-term dependencies. In the absence of the regularization term in the learning equation, it appears that the network takes a longer time to adapt. Furthermore, the network struggles to re-adapt when there is a major change in input patterns. When it comes to the developed hardware model, which uses regularization, a similar trend as the software model has been observed. However, one may notice there is a gap in performance ($\sim$3.9%) between the software and hardware model, which can attributed to several reasons. Among these are the non-idealities of memristor devices, limited precision of the ADC used when modulating the conductance of the memristors, and undesired leakage in the sample and hold circuit leading to a drop in the stored charges.
Table 2. The forecasting wMAPE of univariate stationary and non-stationary benchmarks using the proposed memristor-based ESN when trained with LMS, LMS with L2 regularization, and when using point and leaky-integrated neurons (LINs) in the reservoir layer.
Benchmark Forecast LMS (LIN) LMS+L2 LMS+L2 (LIN)
PJM-Energy 50-Step 0.073$\pm$0.0011 0.087$\pm$0.0024 0.061$\pm$0.0084
100-Step 0.075$\pm$0.010 0.092$\pm$0.023 0.066 $\pm$0.0089
Mackey-Glass 50-Step 0.053$\pm$0.0069 0.079$\pm$0.0273 0.047$\pm$0.0004
100-step 0.060$\pm$0.011 0.082$\pm$0.033 0.047$\pm$0.0067
Daily-Temp 50-Step 0.075$\pm$0.0052 0.097$\pm$0.0189 0.073$\pm$0.0015
100-step 0.093$\pm$0.029 0.0105$\pm$0.025 0.083$\pm$0.0172
NARMA10 50-Step 0.191$\pm$0.0039 0.198$\pm$0.0069 0.189$\pm$0.0034
100-Step 0.195$\pm$0.0039 0.211$\pm$0.0192 0.191$\pm$0.0039
Table 2 shows the wMAPE when forecasting data from stationary and non-stationary benchmarks for various time steps. The wMAPE is calculated when training the proposed memristor-based ESN accelerator with LMS, and LMS with L2 regularization using point and leaky-integrated neurons in the reservoir layer. It can be seen that having leaky-integrated neurons (LINs) significantly improves network performance as it enables the ESN to form temporal dependencies with controllable dynamic updates.
@@ 5.2. Device Failure Effect
There are several types of device failure a network has to deal with especially when using memristor devices. The memristor devices may experience a deviation in characteristics such as a change in the device conductance range due to excessive device switching, known as an aging fault, or it may have a fault in the device right after the fabrication process such as stuck-at-fault. In this work, we will primarily focus on stuck-at fault because it is most common and may severely impact network performance (Xu et al., 2018). There are three types of stuck-at fault: stuck-at, stuck-on, and stuck-off. Previous work has shown that stuck-at replicates the memristor device failure after the forming process and such fault has marginal effect on network performance (Zyarah and Kudithipudi, 2019b). Thus, this work solely investigates the impact of stuck-on (high-conductance state) and stuck-off (low-conductance state) faults on predicting future events in time-series data. In Fig. 5, we show the change in wMAPE when using the PJM energy dataset to predict the energy consumption of the load for the next 50 hours in the presence of various levels of stuck-on and stuck-off faults occurring in the readout and reservoir layers. The fault impact is investigated in both 1M1R and 2M crossbar structure. In 1M1R structure, it is found that regardless of the fault type, below 10% its impact can be deemed marginal causing approximately $\pm$1.3change in the wMAPE value. However, this impact exacerbates beyond 10% and seems to have an exponential negative impact on wMAPE, see Fig. 5-(b). In 2M structure, we may observe similar impact, but it can be suppressed by leveraging the programmable nature of memristor devices. When using two memristors to represent each synaptic weight, one can: i) leverage the intact memristors to tune the weights modeled by faulty devices and here the weights will not be frozen but rather have limited dynamic range, ii) enforce sparsity, i.e. zero weight values, via equating the conductance of the intact and faulty devices. Fig. 5-(c) illustrates the wMAPE in the presence of stuck-on and stuck-at faults when we enforce sparsity.
Figure 5. (a) The wMAPE of the developed ESN models (SW and HW) is calculated while predicting the load energy consumption for 50 hours, computed every 250 samples. (b) and (c) show the impact of various levels of stuck-on and stuck-off faults in memristor devices on the network performance while performing time-series forecasting task when using 1M1R and 2M crossbar structures, respectively.